# Splat-MOVER: Multi-Stage, Open-Vocabulary Robotic Manipulation via Editable Gaussian Splatting

### v2

@@ IV Affordance-and-Semantic-Knowledge Gaussian Splatting
In this section, we present ASK-Splat, an Affordance-and-Semantic Knowledge Gaussian Splatting representation, that provides object-level semantic information and grasp affordance, enabling open-set queries in natural language. ASK-Splat consists of: (1) a semantic module, grounding language embeddings in a $3D$ Gaussian Splatting scene, enabling the generation of similarity (or relevancy) maps given open-vocabulary text prompts; and (2) an affordance module, grounding grasp affordance embeddings to provide highly-localized object-specific regions amenable to grasping. We illustrate these components in Figure 1, where we assume that we have a dataset $\mathcal{D}$ of RGB images, given by ${\mathcal{D}=\{\mathcal{I}_{1},\ldots,\mathcal{I}_{N}\}}$. We discuss these components in greater detail in the following subsections.
Figure 1: ASK-Splat is trained with: (1) RGB images of the scene and is initialized with a sparse point cloud from structure-from-motion (in the green box); (2) CLIP latent semantic features computed from an encoder-decoder model (in the blue box); and (3) affordance codes from a vision-affordance foundation model (in the purple box), e.g., VRB.
IV-A Grounding Language Semantics in $3$D Gaussian Splatting
In NeRF-based methods, semantic knowledge is distilled into $3$D via the introduction of a semantic feature field (defined by MLPs), mapping each point in both the free and occupied regions of the scene to a semantic feature embedding. However, we note that learning the semantic information over both free and unoccupied space is generally inefficient, as many human and robotics tasks only require semantic knowledge of occupied space, e.g., when asked to pick up a cup, a human typically focuses on the semantic query ‘cup,’ which lies within the occupied region of the scene. As a result, in ASK-Splat, we ground language embeddings in the occupied regions of the scene, represented by the $3$D Gaussians.
Remark 1 (Computation Efficiency).
A naive approach to encode semantic information within a Gaussian Splatting representation would be to assign an additional parameter ${f\in\mathbb{R}^{C}}$ assigned to each Gaussian, representing the semantic feature embedding associated with the Gaussian. However, such an approach introduces significant memory and computation challenges, effectively eliminating the real-time rendering rates of Gaussian Splatting. Moreover, the memory available on commercially available hardware might be insufficient for representing the scene, especially large scenes with many Gaussians.
@@ IV-A Grounding Language Semantics in $3$D Gaussian Splatting
To preserve the real-time rendering speed of Gaussian Splatting, we learn a lower-dimensional latent space associated with the high-dimensional semantic feature embeddings, and subsequently, leverage the lower-dimensional latent features for semantic knowledge distillation, which we describe in the following discussion. Given an RGB image ${\mathcal{I}\in\mathbb{R}^{H\times W\times 3}}$, we begin by extracting ‘ground-truth’ image embeddings from the vision-language foundation model CLIP, denoted by ${f_{\mathrm{gt}}\in\mathbb{R}^{H_{f}\times W_{f}\times C}}$, using MaskCLIP [4]. We note that although the image embeddings from [4] are sometimes described to be dense, the resulting image embeddings for the original input image are essentially patch-based, with ${H_{f}<H}$ and ${W_{f}<W}$. Further, we note that the dimension of $f_{\mathrm{gt}}$ (denoted by $C$) depends on the CLIP model used, with common values ranging between $512$ (for the CLIP-ViT models) and $1024$ (for the CLIP-ResNet models).
@@
Subsequently, we learn a lower-dimensional encoding of the semantic embeddings using an autoencoder, consisting of an encoder ${g_{\phi}^{\mathrm{enc}}:\mathbb{R}^{H_{f}\times W_{f}\times C}\mapsto\mathbb{R}^{H_{f}\times W_{f}\times l}}$, mapping semantic embeddings of dimension $C$ from the semantic feature space to the semantic latent space, and a decoder ${g_{\theta}^{\mathrm{dec}}:\mathbb{R}^{H_{f}\times W_{f}\times l}\mapsto\mathbb{R}^{H_{f}\times W_{f}\times C}}$, mapping the semantic latent embeddings of dimension $l$ to the semantic feature space, with parameters $\phi$ and $\theta$, respectively. We utilize feedforward MLPs in defining the encoder and decoder (depicted in Figure 1) with ${l=3\ll C}$. However, we note that larger values of $l$ generally result in more expressive semantic scene representations, at the expense of increased memory and rendering costs. We train the autoencoder with the loss function $\mathcal{L}_{g}$, given by:
@@
Given a trained encoder, we map the ground-truth image embeddings from CLIP to the lower-dimensional latent space and distill the lower-dimensional embeddings into the Gaussian Splatting representation. We augment the attributes of each Gaussian with an additional parameter ${f_{s}\in\mathbb{R}^{l}}$, denoting the semantic feature of the Gaussian. To render $2D$ semantic feature maps of the scene, we utilize the same tile-based rasterization procedure presented in [17], culling $3$D Gaussians whose $99\%$ confidence ellipsoid do not intersect the view frustum associated with the pose of the camera. We optimize the semantic feature parameters using the loss function:
@@
IV-B Grounding Affordance in $3$D Gaussian Splatting
ASK-Splat embeds object-specific grasp affordances directly within the $3$D Gaussian Splatting environment, enabling real-time queries of relevant regions of an object for grasping. By distilling visual grasp affordances from vision-affordance foundation models trained on large human-object interaction datasets, such as VRB [6], into the scene representation, we endow robots with the ability to identify the parts of an object that a human is more likely to grasp, capturing common-sense human knowledge and experience in interacting with objects.
In the subsequent discussion, we present our procedure for embedding grasp affordances in ASK-Splat. We leverage the $3$D Gaussian primitives in embedding $2D$ visual grasp affordances in $3D$, without resorting to a neural-based approach. We obtain the ground-truth affordance score by evaluating the training dataset $\mathcal{D}$ using the vision-affordance foundation model. By augmenting the attributes of each Gaussian in the scene with an additional parameter ${\beta\in\left[0,1\right]}$, representing the score of the Gaussian affording the task of grasping, we ground the grasp affordances within the geometric primitives representing the occupied regions of the scene. We enforce the box constraints on $\beta$ using the sigmoid activation function to provide smooth gradients during the training procedure.
During training, we utilize the same tile-based rasterization procedure discussed in Section IV-A in rendering the $2D$ visual grasp affordance of the scene and optimize the grasp affordance parameter $\beta$ using the loss function:
$\mathcal{L}_{\beta}=$ $\kappa_{\beta}\sum_{i=1}^{\lvert\mathcal{D}\rvert}\left\|\hat{\mathcal{I}}_{i}^{\beta}-\mathcal{I}_{i}^{\beta}\right\|_{2}^{2},$ (5)
@@ IV-B Grounding Affordance in $3$D Gaussian Splatting
which represents the MSE loss between the ground-truth $2D$ visual grasp affordance ${\mathcal{I}_{i}^{\beta}\in\mathbb{R}^{H\times W\times 1}}$ and the rendered visual grasp affordance ^${\hat{\mathcal{I}}_{i}^{\beta}\in\mathbb{R}^{H\times W\times 1}}$. We optimize the affordance parameters concurrently with the non-semantic parameters of each Gaussian. From a trained ASK-Splat scene, we can generate dense $2$D visual grasp affordance maps, as well as sparser $3$D visual grasp affordance maps, by directly evaluating the affordance score associated with each Gaussian.
V Scene-Editing-Enabled Gaussian Splatting
We present SEE-Splat, our module for Scene-Editing-Enabled Gaussian Splatting representations. SEE-Splat harnesses the semantic and affordance knowledge in ASK-Splat to identify and localize relevant objects within a scene for insertion, removal, or modification of the object’s visual or spatial properties. SEE-Splat consists of three components: (i) a semantic component utilizing natural-language queries to generate a relevancy map based on semantic similarity; (ii) a Gaussian masking module that generates a sparse point cloud of the relevant objects, comprising of all semantically-relevant Gaussians; and (iii) a scene transformation component that edits the $3$D scene by inserting, removing, or modifying the geometric, spatial, or visual properties of the Gaussians. In Figure 2, we illustrate each component in editing a real-world Kitchen scene to visualize the action of moving the saucepan from the table to the electric stove, showing localization of a saucepan, extraction of the semantically-relevant Gaussians, and transformation of the relevant Gaussians. Together, these components enable robots to visualize the effects of their interactions with other objects in a virtual $3$D scene, prior to executing these actions in the real-world. Essentially, SEE-Splat provides a digital twin of the scene. Moreover, provided sensor feedback is available, SEE-Splat enables real-time visualization of the real-world in a virtual environment, enabling the Gaussian primitives to accurately reflect the real-time geometry and visual properties of objects in the real-world. We discuss each component in the following subsections.
Figure 2: Given a natural-language prompt and a desired scene transformation, SEE-Splat leverages $3$D ASK-Splat for open-vocabulary scene-editing via semantic localization of relevant Gaussian primitives in the scene, followed by $3$D masking and transformation of these Gaussians. We visualize the Gaussians as as a point cloud, with well-defined geometry.
V-A Semantic Localization via ASK-Splat

### v3

@@ IV Affordance-and-Semantic-Knowledge Gaussian Splatting
Given a set $\mathcal{D}$ of RGB images, given by ${\mathcal{D}=\{\mathcal{I}_{1},\ldots,\mathcal{I}_{N}\}}$, ASK-Splat embeds two kinds of features in the GSplat scene: (1) CLIP features, and (2) grasp affordance features. Using the CLIP features, ASK-SPlat generates a 3D heat map in response to an open-vocabulary text prompt, which denotes the semantic relevance of regions in the scene to the prompt. E.g., the prompt “apple” will give a 3D heat map that is hottest over the apple, but also somewhat hot over the pear and orange (which are semantically “close” to apple). Although the affordance feature field cannot be queried directly in natural language, it provides information on the graspability of all points in the scene. We use the CLIP features to mask out an object of interest, and the affordance features to highlight graspable locations within the mask. We illustrate these components in Figure 1.
Figure 1: ASK-Splat grounds $2$D visual attributes (e.g, color and lighting effects), grasp affordance, and semantic embeddings within a $3$D GSplat representation and is trained entirely from RGB images. Leveraging $3$D ASK-Splat, SEE-Splat enables open-vocabulary scene-editing via semantic localization of relevant Gaussian primitives in the scene, followed by $3$D masking and transformation ${\xi(t)}$ of these Gaussians.
@@
Grounding Language Semantics in 3D GSplats A naive approach to encode semantic information within a Gaussian Splatting representation would be to assign an additional parameter ${f\in\mathbb{R}^{C}}$ assigned to each Gaussian, representing the semantic feature embedding associated with the Gaussian. However, such an approach introduces significant memory and computation challenges, effectively eliminating the real-time rendering rates of Gaussian Splatting. To preserve the real-time rendering speed of Gaussian Splatting, we learn a lower-dimensional latent space associated with the high-dimensional semantic feature embeddings, and subsequently, leverage the lower-dimensional latent features for semantic knowledge distillation. Given an RGB image ${\mathcal{I}\in\mathbb{R}^{H\times W\times 3}}$, we begin by extracting ‘ground-truth’ image embeddings from the vision-language foundation model CLIP, denoted by ${f_{\mathrm{gt}}\in\mathbb{R}^{H_{f}\times W_{f}\times C}}$, using MaskCLIP [13]. We note that although the image embeddings from [13] are sometimes described to be dense, the resulting image embeddings for the original input image are essentially patch-based, with ${H_{f}<H}$ and ${W_{f}<W}$. Further, we note that the dimension of $f_{\mathrm{gt}}$ (denoted by $C$) depends on the CLIP model used, with common values ranging between $512$ (for the CLIP-ViT models) and $1024$ (for the CLIP-ResNet models). Subsequently, we learn a lower-dimensional encoding of the semantic embeddings using an autoencoder, consisting of an encoder ${g_{\phi}^{\mathrm{enc}}:\mathbb{R}^{H_{f}\times W_{f}\times C}\mapsto\mathbb{R}^{H_{f}\times W_{f}\times l}}$, mapping semantic embeddings of dimension $C$ from the semantic feature space to the semantic latent space, and a decoder ${g_{\theta}^{\mathrm{dec}}:\mathbb{R}^{H_{f}\times W_{f}\times l}\mapsto\mathbb{R}^{H_{f}\times W_{f}\times C}}$, mapping the semantic latent embeddings of dimension $l$ to the semantic feature space, with parameters $\phi$ and $\theta$, respectively, with ${l=3\ll C}$, depicted in Figure 1. We distill the latent CLIP features into the Gaussian Splatting representation, by augmenting the attributes of each Gaussian with an additional parameter ${f_{s}\in\mathbb{R}^{l}}$, denoting the semantic feature of the Gaussian, which are optimized via gradient descent. We describe the distillation and optimization procedures in greater detail in Appendix I.
Grounding Affordance in 3D Gaussian Splatting ASK-Splat embeds object-specific grasp affordances directly within the $3$D Gaussian Splatting environment, yielding a heat map over objects in the scene relating to the graspability. We distill visual grasp affordances from a vision-affordance foundation model trained on large human-object interaction datasets, VRB [14], into the scene representation, although we note that other vision-affordance foundation models can also be used. This endows robots with the ability to identify the parts of an object that a human is more likely to grasp, capturing common-sense human knowledge and experience in interacting with objects.
We obtain the ground-truth affordance score by evaluating the training dataset $\mathcal{D}$ using VRB and leverage the $3$D Gaussian primitives in embedding the $2$D visual grasp affordance scores in $3$D. By augmenting the attributes of each Gaussian in the scene with an additional parameter ${\beta\in\left[0,1\right]}$, representing the score of the Gaussian affording the task of grasping, we ground the grasp affordances within the geometric primitives representing the occupied regions of the scene. We enforce the box constraints on $\beta$ using the sigmoid activation function to provide smooth gradients during the training procedure. We provide additional details in Appendix I.
Figure 2: Grasp-Splat generates affordance-aligned grasps using the semantic and grasp-affordance knowledge in ASK-Splat. Qualitatively, the proposed grasps lie in regions where a human is more likely to grasp (e.g., the handle of the saucepot and the center of the fruit) and are more likely to result in stable grasps, when executed by a robot.
V Scene-Editing-Enabled Gaussian Splatting
SEE-Splat consists of three components: (i) a semantic component utilizing natural-language queries to generate a relevancy map based on semantic similarity; (ii) a Gaussian masking module that generates a sparse point cloud of the relevant objects, comprising of all semantically-relevant Gaussians; and (iii) a scene transformation component that edits the $3$D scene by inserting, removing, or modifying the geometric, spatial, or visual properties of the Gaussians. In Figure 1, we illustrate each component of SEE-Splat by editing a real-world cooking scene to visualize the action of moving the saucepot from the table to the electric stove, showing localization of the saucepot, extraction of the semantically-relevant Gaussians, and transformation of the relevant Gaussians. Given a natural language query, we generate a semantic similarity heatmap of the object within ASK-Splat, which is used to mask the Gaussians belonging to the object, given a threshold for semantic relevancy. To edit ASK-Splat, SEE-Splat applies a transform ${\xi:\mathcal{G}_{s}\mapsto\mathcal{G}_{s}}$ to the relevant Gaussians to update their spatial attributes, where $\mathcal{G}_{s}$ represents the space of the Gaussian primitives. Together, these components enable robots to visualize the effects of their interactions with other objects in a virtual $3$D scene, prior to executing these actions in the real-world. Essentially, SEE-Splat provides a digital twin of the scene. Moreover, provided sensor feedback is available, SEE-Splat enables real-time visualization of the real-world in a virtual environment, enabling the Gaussian primitives to accurately reflect the real-time geometry and visual properties of objects in the real-world. We provide a more detailed discussion of each component in Appendix II.
VI Affordance-Aligned Grasp Generation
Grasp-Splat utilizes the dense point cloud of the object generated by SEE-Splat to propose grasp configurations for grasping the object, while harnessing the grasp affordance knowledge in ASK-Splat to identify generally more stable grasp configurations associated with the specified object. The proposed grasps are generated from the point cloud using a deep-learned model GraspNet [3], that estimates $6$D grasp poses for a parallel-jaw gripper from a point cloud of the object, along with estimated grasp-quality scores associated with these grasp poses. We note that the grasps generated by GraspNet are not always ideal. Consequently, Grasp-Splat executes a grasp selection procedure to identify more-promising candidate grasps. We hypothesize that leveraging grasp affordance of each part of the object when generating candidate grasps could be essential to identifying better candidate grasps. As a result, we introduce the grasp metric ${\nu:\mathrm{SE}(3)\mapsto\mathbb{R}}$, which computes the grasp affordance at a specified grasp pose ${X\in\mathrm{SE}(3)}$. Grasp-Splat ranks the candidate grasps generated by GraspNet based on the grasp scores given by $\nu(X)$, leveraging the affordance associated with each grasp pose to identify grasp configurations that are more likely to succeed, depicted in Figure 2. Moreover, since GraspNet does not consider the position of the robot relative to the object, the proposed grasps might require post-processing to account for the relative position of the robot. We elaborate more on our grasp-selection procedure and its application to multi-stage robotic manipulation in Appendix III.
TABLE I: Grasping success rate and the percentage of grasps in the affordance region (AGSR) of each object for LERF-TOGO [5], F3RM [6], and Splat-MOVER across $20$ feasible trials are presented in the first column. In addition, we present the pick and place success rates for the two-stage manipulation task across four scenarios where each scenario has two pick and place stages. From top to bottom: Cooking task, where a robot is asked to move a saucepan to an electric burner, followed by placing a fruit in the saucepan; Chopping task, where a robot is asked to move a knife to a chopping board, followed by placing a fruit next to the knife; Cleaning task, where a robot is asked to move a cleaning spray into a bin, followed by placing a sponge next to the bin; and Workshop task, where a robot is asked to move a power drill to a work mat, followed by moving a wooden block next to the drill. The top-performing stat is shown in bold.
[…]
VII Hardware Evaluations
We compare Splat-MOVER to the existing open-vocabulary robotic manipulation methods LERF-TOGO [5] and F3RM [6] in four tasks: the Cooking task, Chopping task, Cleaning task, and Workshop task, illustrated in Table I, executed on a Kinova robot. We utilize the publicly-available implementation of LERF-TOGO provided by the authors. Since F3RM requires collecting human demonstrations in virtual-reality, which we could not provide in our evaluations, we implement F3RM and utilize GraspNet for grasp generation.
Since the NeRF-based methods LERF-TOGO and F3RM are not amenable to scene-editing, we limit our comparisons to the first stage of the manipulation task, where we consider grasping a saucepan, knife in a knife guard, cleaning spray, and power drill. In the hardware experiments on the robot, we describe a candidate grasp as being feasible if the MoveIt planner successfully finds a plan to execute the grasp. We execute $20$ feasible trials for each object, and in Table I, provide the grasping success rate of each method, as well as the percentage of grasps that lie within the affordance region of the object (AGSR), where the affordance region of each object is defined as follows: the entire handle of the saucepan, the entire knife with the guard on, the region of the cleaning spray excluding its cap, and the middle region of the power drill below its top compartment and above its battery compartment. From Table I, Splat-MOVER achieves the best grasping success rates and the highest percentage of grasps in the affordance region in grasping the saucepan, knife, and cleaning spray. LERF-TOGO attains a perfect success rate in grasping the power drill; however, LERF-TOGO always grasped the top of the drill, outside its affordance region, shown in Appendix IV. In contrast, Splat-MOVER achieves a $95\%$ success rate, grasping the object within its affordance region.
In addition, we evaluate the pick-and-place success rate of all the methods in each manipulation task, where the place success rate is conditioned on the number of successful trials in picking the object. The publicly-available implementation of LERF-TOGO does not support the specification of a place location. Hence, we do not evaluate its place success rate. In Table I, we present the pick-and-place success rates in the four tasks. In the Cooking task, Splat-MOVER achieves a perfect success rate in picking up the saucepan, with a place success rate of $60\%$. We note that the place step involves placing the saucepan on an electric burner, which is quite challenging, potentially explaining the lower place success rate. Nonetheless, Splat-MOVER outperforms LERF-TOGO and F3RM in the first stage of the task, the only stage of the task that both methods can perform. We present additional results, in addition to details on the experimental setup, in Appendix IV.
VIII Conclusion
We present Splat-MOVER, a robotics stack for multi-stage open-vocabulary robotic manipulation, consisting of three modules: ASK-Splat, SEE-Splat, and Grasp-Splat. ASK-Splat enables semantic and affordance queries via natural-language interaction to identify relevant objects within $3$D scenes, while SEE-Splat provides real-time, dynamic scene editing, enabling visualization of the evolution of the real-world scene due to the robot’s interaction with objects within the scene. Grasp-Splat builds upon these two modules for affordance-aware grasp generation, necessary for effective multi-stage robotic manipulation. We demonstrate the effectiveness of Splat-MOVER in real-world experiments in comparison to two recent baseline methods. Compared to the existing works, Splat-MOVER endows robots with the unique capability for multi-stage, open-vocabulary manipulation with minimal human inputs, by leveraging distilled grasp affordance knowledge and real-time dynamic scene editing, which are essential to multi-stage manipulation.
IX Limitations and Future Work
We present a few limitations of Splat-MOVER and provide directions for future work. ASK-Splat distills grasp affordance knowledge from foundation models into $3$D Gaussian Splatting scenes. We note that the effectiveness of the distilled grasp affordances and the generalization capability of our model is highly dependent on that of the affordance foundation model. For example, when trained using the VRB foundation model, an ASK-Splat scene may not generalize notably well to markedly-different objects outside of those seen in the EPIC-KITCHENS dataset. Future work will examine enhancing the generalization capability of the proposed ASK-Splat module by training and employing diverse cross-environment, cross-task, vision-affordance foundation models, rather than relying on VRB, which is specifically trained for kitchen tasks. Further, in its current form, the distilled visual grasp affordance does not depend on the orientation of the candidate grasp, limiting Grasp-Splat from fully harnessing grasp affordances in identifying better grasp configurations in $\mathrm{SE}(3)$. Future work will consider methods for extending the computation of grasp affordance to $\mathrm{SE}(3)$. Moreover, future work will seek to integrate fast online Gaussian Splatting into ASK-Splat, eliminating the need for a brief scanning phase of the scene prior to training in addition to speeding up the training procedure. Additionally, we will seek to integrate sensor feedback into SEE-Splat to enable closed-loop, real-time scene editing, improving the accuracy of the scene representation. In addition, by closing the scene-editing loop with sensor feedback, we can ensure that our $3$D scene representation updates dynamically according to unexpected real-world events. For example, if the pan slips out of the robot gripper, SEE-Splat will be able to promptly reflect this change and modify the plans of the robot accordingly.
Appendix A Affordance-and-Semantic-Knowledge Gaussian Splatting
Here, we provide additional details of the distilled knowledge in ASK-Splat.
A-A Grounding Language Semantics in $3$D Gaussian Splatting
@@
In this work, we utilize feedforward MLPs in defining the encoder and decoder with ${l=3\ll C}$. However, we note that larger values of $l$ generally result in more expressive semantic scene representations, at the expense of increased memory and rendering costs. We train the autoencoder with the loss function $\mathcal{L}_{g}$, given by:
@@
Given a trained encoder, we map the ground-truth image embeddings from CLIP to the lower-dimensional latent space and distill the lower-dimensional embeddings into the Gaussian Splatting representation. We assign a semantic feature ${f_{s}\in\mathbb{R}^{l}}$ to each Gaussian. To render $2D$ semantic feature maps of the scene, we utilize the same tile-based rasterization procedure presented in [28], culling $3$D Gaussians whose $99\%$ confidence ellipsoid do not intersect the view frustum associated with the pose of the camera. We optimize the semantic feature parameters using the loss function:
@@
A-B Grounding Affordance in $3$D Gaussian Splatting
@@
We embed grasp affordances in ASK-Splat. During training, we utilize the same tile-based rasterization procedure discussed in Section A-A in rendering the $2D$ visual grasp affordance of the scene and optimize the grasp affordance parameter $\beta$ using the loss function: ${\mathcal{L}_{\beta}=\kappa_{\beta}\sum_{i=1}^{\lvert\mathcal{D}\rvert}\left\|\hat{\mathcal{I}}_{i}^{\beta}-\mathcal{I}_{i}^{\beta}\right\|_{2}^{2},}$ which represents the MSE loss between the ground-truth $2D$ visual grasp affordance ${\mathcal{I}_{i}^{\beta}\in\mathbb{R}^{H\times W\times 1}}$ and the rendered visual grasp affordance ^${\hat{\mathcal{I}}_{i}^{\beta}\in\mathbb{R}^{H\times W\times 1}}$. We optimize the affordance parameters concurrently with the non-semantic parameters of each Gaussian. From a trained ASK-Splat scene, we can generate dense $2$D visual grasp affordance maps, as well as sparser $3$D visual grasp affordance maps, by directly evaluating the affordance score associated with each Gaussian.
Appendix B Scene-Editing-Enabled Gaussian Splatting
We present the components that make up SEE-Splat, our module for Scene-Editing-Enabled Gaussian Splatting representations, that enables the identification and localization of relevant objects within a scene for insertion, removal, or modification of the object’s visual or spatial properties.
B-A Semantic Localization via ASK-Splat